OpenStudy (anonymous):

How do i solve this equation? 31=8g-9

OpenStudy (anonymous):

You can do like this: 31 = 8g -9 +9 to both sides 31+9 = 8g -9 +9 40 = 8g Divide both sides by 8: 40/8 = 8g/8 By cancelling the 8 on the right side, and by dividing the 40 by 8 you get the answer: g=5 or 5=g if you like it that way.. :)